# Differentially Private Graph Diffusion with Applications in Personalized PageRanks

### v1

@@ Appendix C Discussion on Personalized Graph Diffusion
since $[\mathbf{s}]_{u}=[\mathbf{s}]_{v}=0$. Thus, when analyzing noisy graph diffusion in a personalized scenario, this distinction results in a better bound for Lemma B.2 and  B.3. Further, Eq. (31) demonstrate that we can relax the thresholding over seed node to seek for better privacy-utility trade-offs. Therefore, for uniform noise scheduling, we obtain the corresponding results in Theorem 4:
@@ D.4 Ablation Study
Noise Distributions: Laplace versus Gaussian. We investigate the effectiveness of injecting Laplace noise as compared to Gaussian noise within our analysis framework. Recall that standard PABI analysis has primarily focused on the case of Gaussian noise. We report the ranking performance in Fig. 9 (Left). Within a strong privacy regime where $\epsilon$ ranges from $0.1$ to $1$, the performance gap between the two noise types is approximately 0.04. This finding supports the superiority of Laplace noise injection for graph diffusion processes in $\ell_{1}$ space and is of practical importance.
@@ Appendix F Proof of Lemmas
From definition that $f,f^{\prime}$ are $1$-Lipschitz continuous, while for $\phi_{k}$ and $\phi_{k}^{\prime}$, since $\phi_{k}(\mathbf{x})=\left(\gamma_{1,k}\mathbf{P}+\gamma_{2,k}\right)\mathbf{x}+\gamma_{3,k}\mathbf{s},\phi_{k}^{\prime}(\mathbf{x})=\left(\gamma_{1,k}\mathbf{P}^{\prime}+\gamma_{2,k}\right)\mathbf{x}+\gamma_{3,k}\mathbf{s}$ where $\gamma_{k,1},\gamma_{k,2},\gamma_{k,3}$ are all nonnegative constants such that $\sum_{j=1}^{3}\gamma_{k,j}=1$, we have
$\|\phi_{k}(\mathbf{x})-\phi_{k}(\mathbf{y})\|_{1}=\|\left(\gamma_{1,k}\mathbf{P}+\gamma_{2,k}\right)\mathbf{x}+\gamma_{3,k}\mathbf{s}-\left(\gamma_{1,k}\mathbf{P}+\gamma_{2,k}\right)\mathbf{y}-\gamma_{3,k}\mathbf{s}\|_{1}$ (33)
$\leq$ $\|(\gamma_{1,k}\mathbf{P}+\gamma_{2,k})(\mathbf{x}-\mathbf{y})\|_{1}\leq\gamma_{1,k}\|\mathbf{P}(\mathbf{x}-\mathbf{y})\|_{1}+\gamma_{2,k}\|\mathbf{x}-\mathbf{y}\|_{1}$ (34)
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $(\gamma_{1,k}+\gamma_{2,k})\cdot\|\mathbf{x}-\mathbf{y}\|_{1}$ (35)
where $(a)$ stems from the property that the random walk matrix $\mathbf{P}$ is a left stochastic matrix with each column summing to 1, thus constituting a $1$-contraction mapping w.r.t $\|\cdot\|_{1}$. Similar derivation can be done for $\phi_{k}^{\prime}$, and we conclude that $\phi_{k},\phi_{k}^{\prime}$ are $(\gamma_{1,k}+\gamma_{2,k})$-contraction mappings.
Now, we will focus on distortion analysis: given any $\mathbf{x}\in\mathbb{R}^{n}$, WLOG, we assume edge sets of two graphs satisfies =$\mathcal{E}=\mathcal{E}^{\prime}\cup\{(1,2)\}$, i.e. we drop edge $(1,2)$ on graph $\mathcal{G}$ which lands in graph $\mathcal{G}^{\prime}$.
$\|\phi_{k}(f(\mathbf{x}))-\phi_{k}^{\prime}(f^{\prime}(\mathbf{x}))\|_{1}$ (36)
$\leq$ $\|\phi_{k}(f(\mathbf{x}))-\phi_{k}^{\prime}(f(\mathbf{x}))\|_{1}+\|\phi_{k}^{\prime}(f(\mathbf{x}))-\phi_{k}^{\prime}(f^{\prime}(\mathbf{x}))\|_{1}$ (37)
$=$ $\|(\gamma_{1,k}\mathbf{P}+\gamma_{2,k}\mathbf{I})f(\mathbf{x})-(\gamma_{1,k}\mathbf{P}^{\prime}+\gamma_{2,k}\mathbf{I})f(\mathbf{x})\|_{1}$
$+\|(\gamma_{1,k}\mathbf{P}^{\prime}+\gamma_{2,k}\mathbf{I})f(\mathbf{x})-(\gamma_{1,k}\mathbf{P}^{\prime}+\gamma_{2,k}\mathbf{I})f^{\prime}(\mathbf{x})\|_{1}$ (38)
$\leq$ $\gamma_{1,k}\|(\mathbf{P}-\mathbf{P}^{\prime})f(\mathbf{x})\|_{1}+(\gamma_{1,k}+\gamma_{2,k})\cdot\|f(\mathbf{x})-f^{\prime}(\mathbf{x})\|_{1}$ (39)
$=$ $\gamma_{1,k}\|(\mathbf{A}\mathbf{D}^{-1}-\mathbf{A}^{\prime}\mathbf{D}^{\prime-1})f(\mathbf{x})\|_{1}+(\gamma_{1,k}+\gamma_{2,k})\cdot\|f(\mathbf{x})-f^{\prime}(\mathbf{x})\|_{1}$ (40)
Further, we consider $\mathbf{A}\mathbf{D}^{-1}-\mathbf{A}^{\prime}\mathbf{D}^{\prime-1}$, define $n=|\mathcal{V}|$:
$\mathbf{A}\mathbf{D}^{-1}-\mathbf{A}^{\prime}\mathbf{D}^{\prime-1}=$ $\begin{bmatrix}0&\frac{1}{d_{2}}&\frac{a_{13}}{d_{3}}&\cdots&\frac{a_{1n}}{d_{n}}\\ \frac{1}{d_{1}}&0&\frac{a_{23}}{d_{3}}&\cdots&\frac{a_{2n}}{d_{n}}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \frac{a_{n1}}{d_{1}}&\frac{a_{n2}}{d_{2}}&\frac{a_{n3}}{d_{3}}&\cdots&\frac{a_{nn}}{d_{n}}\end{bmatrix}-\begin{bmatrix}0&0&\frac{a_{13}}{d_{3}}&\cdots&\frac{a_{1n}}{d_{n}}\\ 0&0&\frac{a_{23}}{d_{3}}&\cdots&\frac{a_{2n}}{d_{n}}\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ \frac{a_{n1}}{d_{1}-1}&\frac{a_{n2}}{d_{2}-1}&\frac{a_{n3}}{d_{3}}&\cdots&\frac{a_{nn}}{d_{n}}\end{bmatrix}$ (41)
$=$ $\begin{bmatrix}0&\frac{1}{d_{2}}&0&\cdots&0\\ \frac{1}{d_{1}}&0&0&\cdots&0\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ a_{n1}(\frac{1}{d_{1}}-\frac{1}{d_{1}-1})&a_{n2}(\frac{1}{d_{2}}-\frac{1}{d_{2}-1})&0&\cdots&0\end{bmatrix}$ (42)
Therefore, we have
$\|(\mathbf{A}\mathbf{D}^{-1}-\mathbf{A}^{\prime}\mathbf{D}^{\prime-1})f(\mathbf{x})\|_{1}$ (43)
$=$ $\frac{|[f(\mathbf{x})]_{2}|}{d_{2}}+\frac{|[f(\mathbf{x})]_{1}|}{d_{1}}+\sum_{j\in\mathcal{N}(1)\backslash\{2\}}\frac{|[f(\mathbf{x})]_{1}|}{d_{1}(d_{1}-1)}+\sum_{l\in\mathcal{N}(2)\backslash\{1\}}\frac{|[f(\mathbf{x})]_{2}|}{d_{2}(d_{2}-1)}$ (44)
$=$ $2\left(\frac{|[f(\mathbf{x})]_{1}|}{d_{1}}+\frac{|[f(\mathbf{x})]_{2}|}{d_{2}}\right)\leq 4\max\left\{\frac{|[f(\mathbf{x})]_{1}|}{d_{1}},\frac{|[f(\mathbf{x})]_{2}|}{d_{2}}\right\}\leq 4\max_{\mathbf{x}\in\mathbb{R}^{n},i}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|$ (45)
where $\mathcal{N}(i)$ denote the neighbors of node $i$. Similarly, the above derivation works for $\mathcal{E}\cup\{(1,2)\}=\mathcal{E}^{\prime}$ with same bound by adding and subtracting $(\gamma_{1,k}\mathbf{P}+\gamma_{2,k}\mathbf{I})f^{\prime}(\mathbf{x})$ in Eq. (36). Since $(1,2)$ can be replaced by an arbitrary edge, $\max_{\mathbf{x}\in\mathbb{R}^{n},i}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|$ upper bounds any edge-level adjacency graphs. Therefore, define $\gamma_{\text{max}}=\max_{j\in\{1,2\},k\in[K]}|\gamma_{j,k}|,\gamma_{\text{max}}^{(1)}=\max_{k\in[K]}|\gamma_{1,k}|$,
@@
$\|\phi_{k}(f(\mathbf{x}))-\phi_{k}^{\prime}(f^{\prime}(\mathbf{x}))\|_{1}\leq$ $4\gamma_{1,k}\max_{\mathbf{x}\in\mathbb{R}^{n},i}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|+(\gamma_{1,k}+\gamma_{2,k})\Delta f$ (46)
$\leq$ $4\gamma_{\text{max}}^{(1)}\cdot\max_{\mathbf{x}\in\mathbb{R}^{n}}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|+\gamma_{\text{max}}\cdot\Delta f$ (47)
where $\Delta f=\sup_{\mathbf{x}\in\mathbb{R}^{n}}\|f(\mathbf{x})-f^{\prime}(\mathbf{x})\|_{1}$ defines the sensitivity of graph-dependent thresholding function.
@@
WLOG, we assume edge sets of two graphs satisfies =$\mathcal{E}=\mathcal{E}^{\prime}\cup\{(1,2)\}$. During no randomness phase, we have
[…]
Since two diffusion process start from the same seed $\mathbf{s}$, from Lemma B.1, we have
$\|\mathbf{s}_{1}-\mathbf{s}_{1}^{\prime}\|_{1}=$ $\|\left(\gamma_{1,1}\mathbf{P}+\gamma_{2,1}\right)f(\mathbf{s})+\gamma_{3,1}\mathbf{s}-\left(\gamma_{1,1}\mathbf{P}^{\prime}+\gamma_{2,1}\right)f^{\prime}(\mathbf{s})-\gamma_{3,1}\mathbf{s}\|_{1}$ (49)
$\stackrel{{\scriptstyle(a)}}{{\leq}}$ $4\gamma_{\text{max}}^{(1)}\cdot\left|\frac{[f(\mathbf{s})]_{i}}{d_{i}}\right|+\gamma_{\text{max}}\cdot\Delta f\stackrel{{\scriptstyle(b)}}{{=}}(4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}})\eta$ (50)
where $(a)$ is derived from Lemma B.1 and $(b)$ is from the definition of $f$. The bound is tight when the differed edge is attached to the nodes with lowest degrees, which demonstrate that adding or removing an edge attached two low-degree nodes will lead to significant privacy leakage.
For any step $l\geq 1$, we have
$\|\mathbf{s}_{l}-\mathbf{s}_{l}^{\prime}\|_{1}=\|\phi_{l}(f(\mathbf{s}_{l-1}))-\phi_{l}^{\prime}(f^{\prime}(\mathbf{s}_{l-1}^{\prime}))\|_{1}$ (51)
$\leq$ $\|\phi_{l}(f(\mathbf{s}_{l-1}))-\phi_{l}^{\prime}(f^{\prime}(\mathbf{s}_{l-1}))\|_{1}+\|\phi_{l}^{\prime}(f^{\prime}(\mathbf{s}_{l-1}))-\phi_{l}^{\prime}(f^{\prime}(\mathbf{s}_{l-1}^{\prime}))\|_{1}$ (52)
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $(4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}})\eta+\gamma_{\text{max}}\|\mathbf{s}_{l}-\mathbf{s}_{l}^{\prime}\|_{1}$ (53)
where $(a)$ arises from Lemma B.1 and the Lipschitzness of $\phi_{l}\circ f$. Define $\rho_{\text{diff}}=(4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}})\eta$, thus, by induction, we conclude that
$\|\mathbf{s}_{m}-\mathbf{s}_{m}^{\prime}\|_{1}\leq$ $\gamma_{\text{max}}\cdot\|\mathbf{s}_{m-1}-\mathbf{s}_{m-1}^{\prime}\|_{1}+\rho_{\text{diff}}$ (54)
$\leq$ $\gamma_{\text{max}}^{2}\cdot\|\mathbf{s}_{m-2}-\mathbf{s}_{m-2}^{\prime}\|_{1}+\rho_{\text{diff}}(1+\gamma_{\text{max}})$ (55)
$\leq$ $\cdots\leq\gamma_{\text{max}}^{m-1}\cdot\|\mathbf{s}_{1}-\mathbf{s}_{1}^{\prime}\|_{1}+\rho_{\text{diff}}(1+\gamma_{\text{max}}+\cdots+\gamma_{\text{max}}^{m-2})$ (56)
$\leq$ $\frac{(1-\gamma_{\text{max}}^{m})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}$ (57)
F.3 Proof of Lemma B.3
@@
Lemma Proof. With the above result, we can upper bound the Rényi divergence over joint noise distributions. We consider two cases for the chosen parameter $\tau$:
(1) When $\tau>m$,
@@
where $(a)$ arises from strong composition rule for Rényi divergence (Lemma A.4), $(b)$ is from the definition of $\xi_{k}^{(2)},\tilde{\xi}_{k}^{(2)}$ with shift $\mathbf{h}_{k}=(\phi_{k}\circ f)(\mathbf{x})-(\phi_{k}^{\prime}\circ f^{\prime})(\mathbf{x})$, and $(c)$ is derived by combining the results that the $\ell_{1}$ norm of the shift is upper bounded by $\|\mathbf{h}_{k}\|_{1}\leq\rho_{\text{diff}}=4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}}$ (Lemma B.1) and the above Laplace bound under Rényi divergence (Eq. (58)) with $g_{\alpha}(\sigma,\rho)=\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}\rho)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}\rho_{\text{diff}}))$.
(2) When $\tau=m$, the shift between noise terms $\xi_{m+1}^{(2)}$ and $\tilde{\xi}_{m+1}^{\prime(2)}$ differs from $\rho_{\text{diff}}$. Instead, it should account for the shift of the entire noise-free diffusion and the distortion incurred at step $m+1$. Specifically, define $\lambda_{m+1}=\frac{(1-\gamma_{\text{max}}^{m+1})\rho_{\text{diff}}}{(1-\gamma_{\text{max}})}$ (Lemma B.2, and the distortion at step $m+1$). Therefore, following the above analysis, we have
$\mathcal{D}_{\alpha}(\xi_{m+1:K}^{(2)}\|\tilde{\xi}_{m+1:K}^{(2)}])\leq$ $\sum_{k=m+1}^{K}\sup_{\zeta_{m+1:k-1}}\mathcal{D}_{\alpha}(\xi_{k}^{(2)}|\xi_{m+1:k-1}^{(2)}=\zeta_{m+1:k-1}\|\tilde{\xi}_{k}^{\prime(2)}|\tilde{\xi}_{m+1:k-1}^{\prime(2)}=\zeta_{m+1:k-1})$ (61)
$\leq$ $g_{\alpha}(\sigma_{m+1},\lambda_{m+1})+\sum_{k=m+2}^{K}g_{\alpha}(\sigma_{k},\rho_{\text{diff}})$ (62)
Summarizing the above, we prove the lemma.
Proof of Result. Now consider Laplacian distribution, for $\mathbf{h}\in\mathbb{R}^{|\mathcal{V}|}$,
$\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i=1}^{|\mathcal{V}|}\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}|\mathbf{h}_{i}|)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma})|\mathbf{h}_{i}|)$ (63)
@@
$\mathop{\text{maximize}}\limits_{h}\sum_{i=1}^{|\mathcal{V}|}\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}|\mathbf{h}_{i}|)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma})|\mathbf{h}_{i}|)$ (64)

### v4

@@ Appendix C Discussion on Personalized Graph Diffusion
since $[\mathbf{s}]_{u}=[\mathbf{s}]_{v}=0$. Thus, when analyzing noisy graph diffusion in a personalized scenario, this distinction results in a better bound for Lemma B.2. Further, Eq. (29) demonstrate that we can relax the thresholding over seed node to seek for better privacy-utility trade-offs. Therefore, for uniform noise scheduling, we obtain the corresponding results in Theorem 4:
@@ D.4 Ablation Study
Noise Distributions: Laplace versus Gaussian. We investigate the effectiveness of injecting Laplace noise as compared to Gaussian noise within our analysis framework. Recall that standard PABI analysis has primarily focused on the case of Gaussian noise. We report the ranking performance in Fig. 10 (Left). Within a strong privacy regime where $\epsilon$ ranges from $0.1$ to $1$, the performance gap between the two noise types is approximately $0.05$. This finding supports the superiority of Laplace noise injection for graph diffusion processes in $\ell_{1}$ space and is of practical importance.
@@ Appendix F Proof of Lemmas
Given any $\mathbf{x}\in\mathbb{R}^{n}$, without loss of generality, we assume that the edge sets of two graphs satisfy $\mathcal{E}=\mathcal{E}^{\prime}\cup{(1,2)}$, i.e., we remove edge $(1,2)$ from graph $\mathcal{G}$, resulting in graph $\mathcal{G}^{\prime}$. Furthermore, let $\mathcal{N}(i)$ denote the neighbors of node $i$, and define the following index sets: $A:=\mathcal{N}(1)\backslash\{2\}$, $B:=\mathcal{N}(2)\backslash\{1\}$, $C:=A\cap B$, $A^{\prime}:=A\backslash C$, and $B^{\prime}:=B\backslash C$. Additionally, we can rearrange the node indices in $\mathcal{G}$ such that $A^{\prime}=\{3,\dots,|A^{\prime}|+2\}$, $B^{\prime}=\{|A^{\prime}|+3,\dots,|A^{\prime}|+|B^{\prime}|+2\}$, and $C=\{|A^{\prime}|+|B^{\prime}|+3,\dots,|A^{\prime}|+|B^{\prime}|+|C|+2\}$. We adopt the shorthand notation $f_{i}=[f(\mathbf{x})]_{i}$ and $\Delta f_{i}=[f(\mathbf{x})]_{i}-[f^{\prime}(\mathbf{x})]_{i}$. We first consider the case where $d_{1},d_{2}\geq 2$, and define $m_{1}:=-\frac{f_{1}}{d_{1}(d_{1}-1)},m_{2}:=-\frac{f_{2}}{d_{2}(d_{2}-1)},\Delta m_{1}:=\frac{\Delta f_{1}}{d_{1}-1},\Delta m_{2}:=\frac{\Delta f_{2}}{d_{2}-1}$. Since degree-based thresholding function $f$ is symmetric, WLOG, we can assume $f$ is non-negative, i.e., $f(\mathbf{x})=\min(\max(\mathbf{x},\mathbf{0}),\eta\cdot\mathbf{d})$. We have the following
$\|\phi_{k}(f(\mathbf{x}))-\phi_{k}^{\prime}(f^{\prime}(\mathbf{x}))\|_{1}=\|\phi_{k}(f(\mathbf{x}))-\phi_{k}^{\prime}(f(\mathbf{x}))+\phi_{k}^{\prime}(f(\mathbf{x}))-\phi_{k}^{\prime}(f^{\prime}(\mathbf{x}))\|_{1}$ (31)
$=$ $\|\gamma_{1,k}(\mathbf{P}-\mathbf{P}^{\prime})f(\mathbf{x})+(\gamma_{1,k}\mathbf{P}^{\prime}+\gamma_{2,k}\mathbf{I})(f(\mathbf{x})-f^{\prime}(\mathbf{x}))\|_{1}$ (32)
$=$ $\|\gamma_{1,k}(\mathbf{A}\mathbf{D}^{-1}-\mathbf{A}^{\prime}\mathbf{D}^{\prime-1})f(\mathbf{x})+(\gamma_{1,k}\mathbf{A}^{\prime}\mathbf{D}^{\prime-1}+\gamma_{2,k}\mathbf{I})(f(\mathbf{x})-f^{\prime}(\mathbf{x}))\|_{1}$ (33)
$=$ $\left\|\gamma_{1,k}\begin{bmatrix}0&\frac{1}{d_{2}}&0&\cdots&0\\ \frac{1}{d_{1}}&0&0&\cdots&0\\ \cdots&\cdots&\cdots&\cdots&\cdots\\ a_{n1}(\frac{1}{d_{1}}-\frac{1}{d_{1}-1})&a_{n2}(\frac{1}{d_{2}}-\frac{1}{d_{2}-1})&0&\cdots&0\end{bmatrix}\begin{bmatrix}f_{1}\\ f_{2}\\ \cdots\\ f_{n}\end{bmatrix}\right.$
@@
$\left.+(\gamma_{1,k}\mathbf{A}^{\prime}\mathbf{D}^{\prime-1}+\gamma_{2,k})\begin{bmatrix}\Delta f_{1}\\ \Delta f_{2}\\ \cdots\\ \Delta f_{n}\end{bmatrix}\right\|_{1}$ (34)
$=$ $\|\gamma_{1,k}\cdot[\frac{f_{2}}{d_{2}},\frac{f_{1}}{d_{1}},\underbrace{m_{1},\cdots,m_{1}}_{|A^{\prime}|\text{ Times}},\underbrace{m_{2},\cdots,m_{2}}_{|B^{\prime}|\text{ Times}},\underbrace{m_{1}+m_{2},\cdots,m_{1}+m_{2}}_{|C|\text{ Times}},0,...,0]^{T}$
$+\gamma_{1,k}\cdot[0,0,\underbrace{\Delta m_{1},...,\Delta m_{1}}_{|A^{\prime}|\text{ Times}},\underbrace{\Delta m_{2},...,\Delta m_{2}}_{|B^{\prime}|\text{ Times}},\underbrace{\Delta m_{1}+\Delta m_{2},...,\Delta m_{1}+\Delta m_{2}}_{|C|\text{ Times}}]^{T}+\gamma_{2,k}\cdot[\Delta f_{1},\Delta f_{1},0,...,0]^{T}\|_{1}$ (35)
$\leq$ $|\gamma_{1,k}|\cdot(|\frac{f_{1}}{d_{1}}|+|\frac{f_{2}}{d_{2}}|)+|\gamma_{1,k}|\cdot|A^{\prime}|\cdot|(\frac{1}{d_{1}}-\frac{1}{d_{1}-1})f_{1}+\frac{\Delta f_{1}}{d_{1}-1}|+|\gamma_{1,k}|\cdot|B^{\prime}|\cdot|(\frac{1}{d_{2}}-\frac{1}{d_{2}-1})f_{2}+\frac{\Delta f_{2}}{d_{2}-1}|$
$+|\gamma_{1,k}|\cdot|C|\cdot|(\frac{1}{d_{1}}-\frac{1}{d_{1}-1})f_{1}+\frac{\Delta f_{1}}{d_{1}-1}+(\frac{1}{d_{2}}-\frac{1}{d_{2}-1})f_{2}+\frac{\Delta f_{2}}{d_{2}-1}|+|\gamma_{2,k}|(|\Delta f_{1}|+|\Delta f_{2}|)$ (36)
=$\stackrel{{\scriptstyle(a)}}{{=}}$ $|\gamma_{1,k}|\cdot(\frac{f_{1}}{d_{1}}+\frac{f_{2}}{d_{2}})+|\gamma_{1,k}|\cdot|A^{\prime}|\cdot\frac{|d_{1}\Delta f_{1}-f_{1}|}{d_{1}(d_{1}-1)}+|\gamma_{1,k}|\cdot|B^{\prime}|\cdot\frac{|d_{2}\Delta f_{2}-f_{2}|}{d_{2}(d_{2}-1)}$
$+|\gamma_{1,k}|\cdot|C|\cdot|\frac{d_{1}\Delta f_{1}-f_{1}}{d_{1}(d_{1}-1)}+\frac{d_{2}\Delta f_{2}-f_{2}}{d_{2}(d_{2}-1)}|+|\gamma_{2,k}|\cdot(\Delta f_{1}+\Delta f_{2})$ (37)
$\stackrel{{\scriptstyle(b)}}{{\leq}}$ $|\gamma_{1,k}|\cdot(\frac{f_{1}}{d_{1}}+\frac{f_{2}}{d_{2}})+|\gamma_{1,k}|\cdot\frac{|d_{1}\Delta f_{1}-f_{1}|}{d_{1}}+|\gamma_{1,k}|\cdot\frac{|d_{2}\Delta f_{2}-f_{2}|}{d_{2}}+|\gamma_{2,k}|\cdot(\Delta f_{1}+\Delta f_{2})$ (38)
=$\stackrel{{\scriptstyle(c)}}{{=}}$ $|\gamma_{1,k}|\cdot(\frac{f_{1}}{d_{1}}+\frac{f_{2}}{d_{2}})+|\gamma_{1,k}|\cdot\frac{f_{1}-d_{1}\Delta f_{1}}{d_{1}}+|\gamma_{1,k}|\cdot\frac{f_{2}-d_{2}\Delta f_{2}}{d_{2}}+|\gamma_{2,k}|\cdot(\Delta f_{1}+\Delta f_{2})$ (39)
$=$ $\sum_{j=1}^{2}|\gamma_{1,k}|(f_{j}^{\prime}-\frac{(d_{j}-2)f_{j}}{d_{j}})+|\gamma_{2,k}|\cdot(f_{j}-f_{j}^{\prime})$ (40)
$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\max(4|\gamma_{1,k}|,2(|\gamma_{1,k}|+|\gamma_{2,k}|))\cdot\eta\leq\max(4\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})\cdot\eta$ (41)
where (a) follows from the non-negativity of the function $f$, and (b) is derived from $|A^{\prime}\cap C|=d_{1}-1$ and $|B^{\prime}\cap C|=d_{2}-1$. Equality is obtained when $C=\emptyset$, i.e., when there are no common neighbors between nodes 1 and 2. $(c)$ is obtained from Result 1 and $(d)$ originates from Result 2.
Result 1. Consistent with the above notations, we have $f_{j}-d_{j}\Delta f_{j}\geq 0$ for $j\in\{1,2\}$.
Results 2. Consistent with the above notations, we have $|\gamma_{1,k}|(f_{j}^{\prime}-\frac{(d_{j}-2)f_{j}}{d_{j}})+|\gamma_{2,k}|(f_{j}-f_{j}^{\prime})\leq\max(2|\gamma_{1,k}|,|\gamma_{1,k}|+|\gamma_{2,k}|)\eta$ for $j\in\{1,2\}$.
Proof of Result 1. From the above definitions, we have $f_{j}=[f(\mathbf{x})]_{j}=\min(\max(x_{j},0),\eta\cdot d_{j})$ and $f_{j}-d_{j}\Delta f_{j}=d_{j}f_{j}^{\prime}-(d_{j}-1)f_{j}$ where $x_{j}$ is the $j$-th entry of $\mathbf{x}$. Now, consider the following cases:
  • When $x_{j}\leq 0$, we have $f_{j}-d_{j}\Delta f_{j}=0$.
  • When $x_{j}\in(0,(d_{j}-1)\eta]$, it follows that $f_{j}-d_{j}\Delta f_{j}=d_{j}\eta x_{j}-(d_{j}-1)\eta x_{j}=\eta x_{j}\geq 0$.
  • When $x_{j}\in((d_{j}-1)\eta,d_{j}\eta]$, we obtain $f_{j}-d_{j}\Delta f_{j}=(d_{j}-1)(d_{j}-x_{j})\eta^{2}\geq 0$.
  • When $x_{j}>d_{j}$, $f_{j}-d_{j}\Delta f_{j}=0$.
In conclusion, based on the above cases, the result is proven.
Proof of Result 2. We consider different cases:
  • When $x_{j}\leq 0$, $|\gamma_{1,k}|(f_{j}^{\prime}-\frac{(d_{j}-2)f_{j}}{d_{j}})+|\gamma_{2,k}|(f_{j}-f_{j}^{\prime})=0$.
  • When $x_{j}\in(0,(d_{j}-1)\eta]$, we have
    $|\gamma_{1,k}|(f_{j}^{\prime}-\frac{(d_{j}-2)f_{j}}{d_{j}})+|\gamma_{2,k}|(f_{j}-f_{j}^{\prime})=\frac{2|\gamma_{1,k}|\eta[\mathbf{x}]_{j}}{d_{j}}\leq\frac{2|\gamma_{1,k}|(d_{j}-1)\eta}{d_{j}}$ (42)
  • When $x_{j}\in((d_{j}-1)\eta,d_{j}\eta]$, we obtain
    $|\gamma_{1,k}|(f_{j}^{\prime}-\frac{(d_{j}-2)f_{j}}{d_{j}})+|\gamma_{2,k}|(f_{j}-f_{j}^{\prime})$ (43)
    $=$ $|\gamma_{1,k}|\eta\left(\frac{(d_{j}-1)d_{j}-(d_{j}-2)x_{j}}{d_{j}}\right)+|\gamma_{2,k}|\eta(x_{j}-(d_{j}-1))$ (44)
    $\leq$ $\max(\frac{2|\gamma_{1,k}|(d_{j}-1)\eta}{d_{j}},(|\gamma_{1,k}|+|\gamma_{2,k}|)\eta)\leq\max(2|\gamma_{1,k}|,|\gamma_{1,k}|+|\gamma_{2,k}|)\eta$ (45)
Note that, based on the derivations in $(b)$ and $(d)$, we conclude that the bound is asymptotically tight, with the worst-case scenario occurring when the nodes connected by the perturbed edge have no common neighbors, and their degrees increase to infinity.
Additionally, when $d_{1}=1$ or $d_{2}=1$, it can be directly shown that the distortion is upper bounded by $\max(\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})\cdot\eta$. Note that if we further incorporate Euclidean projection onto $\ell_{1}$ ball, we can regard $\mathscr{P}_{\mathcal{B}}(\mathbf{x})$ as a single input without altering the overall bound.
@@
Lemma Proof. With the above result, we can upper bound the Rényi divergence over joint noise distributions. For $\tau\geq 0$,
@@
where $(a)$ arises from strong composition rule for Rényi divergence (Lemma A.4), $(b)$ is from the definition of $\xi_{k}^{(2)},\tilde{\xi}_{k}^{(2)}$ with shift $\mathbf{h}_{k}=(\phi_{k}\circ f)(\mathbf{x})-(\phi_{k}^{\prime}\circ f^{\prime})(\mathbf{x})$, and $(c)$ is derived by combining the results that the $\ell_{1}$ norm of the shift is upper bounded by $\|\mathbf{h}_{k}\|_{1}\leq\rho_{\text{diff}}=\max(4\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})$ (Lemma B.1) and the above Laplace bound under Rényi divergence (Eq. (46)) with $g_{\alpha}(\sigma,\rho)=\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}\rho)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}\rho_{\text{diff}}))$.
Summarizing the above, we prove the lemma. Note that further considering a projection operator does not affect the distortion, as established in Lemma B.1, and thus leaves the bound derived here unchanged.
Proof of Result. Now consider Laplacian distribution, for $\mathbf{h}\in\mathbb{R}^{|\mathcal{V}|}$, let $h_{i}$ denote the $i$-th entry of the vector, we have
$\mathcal{D}_{\alpha}(\mathcal{L}(\bm{0},\sigma)\|\mathcal{L}(\mathbf{h},\sigma))\stackrel{{\scriptstyle(a)}}{{=}}\sum_{i=1}^{|\mathcal{V}|}\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}|h_{i}|)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}|h_{i}|))$ (49)
@@
$\mathop{\text{maximize}}\limits_{\mathbf{h}}\sum_{i=1}^{|\mathcal{V}|}\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}|h_{i}|)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}|h_{i}|))$ (50)